7. If a = i + 3j j k, b = 2i + 4j j 2k and c = =i + 2j + 4k, find a number λ such that

d = a + b + λc is parallel to the yz-plane.



1
Expert's answer
2021-02-01T11:19:16-0500

Here we have vectors "\vec{a}= \text{\^{i}}+3\text{\^{j}}+\text{\^{k}}" , "\vec{b}=2\text{\^{i}}+4\text{\^{j}}+2\text{\^{k}}" and "\vec{c}=\text{\^{i}}+2\text{\^{j}}+4\text{\^{k}}"


So, we calculate "\vec{d}=\vec{a}+\vec{b}+\lambda\vec{c}"



"\vec{d}=\vec{a}+\vec{b}+\lambda\vec{c}\\
=\text{\^{i}}+3\text{\^{j}}+\text{\^{k}}+2\text{\^{i}}+4\text{\^{j}}+2\text{\^{k}}+\lambda\text{\^{i}}+2\lambda\text{\^{j}}+4\lambda\text{\^{k}}\\
=(3+\lambda)\text{\^{i}}+(7+2\lambda)\text{\^{j}}+(3+4\lambda)\text{\^{k}}"



So, we have "\vec{d}=(3+\lambda)\text{\^{i}}+(7+2\lambda)\text{\^{j}}+(3+4\lambda)\text{\^{k}}"


Now, we know that "\vec{d}" parallel to "yz" plane. This means that "\vec{d}" is "\perp" to "x" axis.


So, here we can equate "\vec{d}\cdot(\text{\^{i}})=0" [As the vector representation of "x" plane is "\text{\^{i}}" ]


So,



"\vec{d}\cdot\text{\^{i}}=0\\
\Rightarrow [((3+\lambda)\text{\^{i}}+(7+ 2\lambda)\text{\^{j}}+(3+4\lambda)\text{\^{k}}]\cdot[\text{\^{i}}]=0\\
\Rightarrow 3+\lambda=0\\
\Rightarrow \lambda=-3"

So, value of "\lambda" such that it is parallel to "yz" plane is



"\fcolorbox{black}{aqua}{$\textcolor{black}{\lambda=-3}$}"